Find the ratio in which line segment joining the points (4,8,10) and (6,10,-8)is divided by yz-plane.

Grade:11

2 Answers

Amit
5 years ago
Hi there. Please post the question once only .. I answered the other posted part. Please refer to that only.
shivam Prajapati
5 years ago
Let be the ratio line segment k:1It located on us. So x=0(6k+1×4)/k+1=06k+4=06k=-4K=-2/3. So answer -2:3